# Comprehensive benchmarking of large language models for RNA secondary structure prediction

### v1

@@ Abstract
Inspired by the success of large language models (LLM) for DNA and proteins, several LLM for RNA have been developed recently. RNA-LLM uses large datasets of RNA sequences to learn, in a self-supervised way, how to represent each RNA base with a semantically rich numerical vector. This is done under the hypothesis that obtaining high-quality RNA representations can enhance data-costly downstream tasks. Among them, predicting the secondary structure is a fundamental task for uncovering RNA functional mechanisms. In this work we present a comprehensive experimental analysis of several pre-trained RNA-LLM, comparing them for the RNA secondary structure prediction task in an unified deep learning framework. The RNA-LLM were assessed with increasing generalization difficulty on benchmark datasets. Results showed that two LLM clearly outperform the other models, and revealed significant challenges for generalization in low-homology scenarios.
@@ Introduction
Ribonucleic acid (RNA) plays a crucial role in many fundamental biological processes, such as gene expression, cell signaling, and post-transcriptional regulation [20, 56]. As in proteins, RNAs function and interaction with other molecules are deeply related to their structure. For example, determining RNA structure is essential for RNA-based therapeutics such as mRNA vaccines [34]. Among all RNA transcripts, only 5% is responsible for protein coding. At the same time, a very large remaining portion is non-coding RNA (ncRNA) [7], which in many cases adopt specific structures to perform important biological functions [6]. Experimental results show that, to some degree, sequence determines the secondary structure, and thus, their function [33, 16, 52]. Therefore, the long-established and yet unresolved RNA secondary structure prediction is major challenge in computational biology today [4, 5].
@@
RNA-LLM attempt to effectively embed RNA bases using deep representation learning, in particular the one developed in the field of natural language processing [18, 39]. Most RNA-LLM are based on bidirectional encoder representations from transformers (BERT), which was designed to generate context-sensitive distributed test token (word) representations [12]. The relevance of LLM for transfer learning from sequences to downstream tasks have been analyzed in detail for proteins [25, 53]. In the case of RNA, the hypothesis is that LLM representation of nucleotide composition and sequence motifs can help characterize the structure and function of a sequence, analogously to how the meaning of a sentence is determined by the grammatical structure of natural language. Thus, word embedding techniques for natural language have been applied to bases for RNA sequences, obtaining a wide variety of BERT-based architectures trained on different databases. However, in spite that some theoretical reviews on the applications and utility of LLM in bioinformatics have recently appeared [26, 50, 35], to the best of our knowledge all the available RNA-LLM have not yet been comparatively evaluated for the secondary structure prediction task, considering a fair experimental setup, with the same datasets and structure prediction model. Since LLM are pre-trained models, it is important to analyze how they were trained to assess generalization capabilities on new sequences/structures, considering homology-aware partitions and cross-family predictions.
@@
Multiple sequence alignment (MSA) inspired by AlphaFold2 [21]. Homology search and MSM of training sequences with RNAcmap3 database [9] as augmented input. Two modules: embedding and MSA transformer. Initial embedding layer and two learnable position-embedding layers. Stack of MSA transformer blocks with a residue and sequence attention layer followed by a feed-forward layer with GELU activation functions. Cross-entropy loss function.
@@ Results
Self-supervised RNA-LLM and prediction model
In the last three years, a number of RNA-LLM have appeared in literature. We summarize in Table 1 the most recent RNA-LLM, included in this study according to their availability as open source tools. All models and pre-training methods are detailed in the Methods section.
For the comparative analysis, embeddings of each RNA-LLM feed the same deep learning architecture (Figure 1a-e) for RNA secondary structure prediction. The rationale behind this is that it should not be necessary to optimize the classifier hyperparameters for each LLM, since they have complex architectures and have already been trained with a large amount of data. Thus, according to the transfer learning paradigm, the embeddings will have enough information to solve the downstream task. Moreover, any difference measured in performance will be due to the LLM and not because of the classifier architecture. This is the only trainable part of the experimental setup, which uses the train-test partitions defined for each benchmark dataset (detailed in Methods).
Figure 1: |||| RNA-LLM embeddings and deep neural architecture of the prediction model. a, Flow diagram for RNA secondary structure predictions. Each LLM was downloaded from its official repository, frozen and used to get per-nucleotide embeddings for each sequence in the benchmark datasets. The embeddings go through a fully connected layer, outer concatenation, two 2D ResNet blocks and a final 2D convolution. This flow is explained in detail in aligned panels b-e. b, One-hot encoding, of size $4\times L$, and LLM representation with size $d\times L$, where $d$ is the embedding dimension and $L$ is the padded sequence length. A fully connected layer reduces the input dimension to $M/2$. c, The $M/2\times L$ projection is transformed to a $M\times L\times L$ tensor using outer concatenation. The position $(i,j)$ contains the concatenated representation of nucleotides $i$ and $j$. d, Then, there are two 2D ResNet blocks with 2D Convolution, Instance Normalization and ReLU. ResNet Blocks 1 has kernel size 1 and ResNet Blocks 2 has kernel size 3. e, A final convolution yields the output scores as a $L\times L$ connection matrix. f-l, RNA-LLM embeddings projected with UMAP for dimensionality reduction. The RNA families of the ArchiveII dataset are highlighted with different colors. Sequences within the same RNA family are expected to be close in the dimensionally reduced space.
@@
Figure 2 shows the comparative results among the RNA-LLM here reviewed in terms of $F_{1}$ violin plots. In each plot the results are presented in order, from best (left) to worst (right) median $F_{1}$. Figure 2a shows the results for 5-fold random partitions on the ArchiveII dataset. It can be seen that most predictors achieve a median performance above $F_{1}=0.60$, except only for the one based on one-hot encoding ($F_{1}=0.57$). The top-3 RNA-LLM achieve a very high performance ($F_{1}>0.90$). Those are ERNIE-RNA and RiNALMo, both with $F_{1}=0.95$, and RNA-FM with $F_{1}=0.91$. RNAErnie and RNA-MSM achieve an intermediate performance, $F_{1}=0.76$ and $F_{1}=0.74$, respectively. Finally, RNABERT is the RNA-LLM with the lowest average performance ($F_{1}=0.62$). For this random $k$-folding, all RNA-LLM except RNABERT achieve higher performance than the classical prediction method (LinearPartition-C [59]).This result shows that the neural architecture used in the classifier has the capability of obtaining high performance in predictions. However, it is well-known that the sequence homology between train and test at random partitions favors methods based on deep learning; thus performance tends to be overly optimistic.
@@
Figure 2: |||| Comparative results among RNA-LLM on the RNA secondary structure prediction task for different benchmark datasets of increasing complexity. Each method has a different color and the performance of LinearPartition-C is shown for reference with a horizontal blue line in all datasets. Next to RNA-LLM performances, sequence length distribution for each dataset is shown in blue. a, ArchiveII 5-fold random cross-validation. A Wilcoxon test for paired samples with Bonferroni correction indicates that all differences are statistically significant ($P<0.0001$, $N=3,864$). b, bpRNA train-test partitions with controlled homology. All differences are statistically significant ($P<0.0001$, $N=1,305$) except for one-hot and RNABERT. c, bpRNA-new dataset, for RNA families not seen during training. All differences are statistically significant ($P<0.0001$, except for RNAErnie versus RNA-MSM and one-hot with $P<0.05$, $N=5,401$). d, PDB-RNA dataset, with RNA sequences extracted from PDB. Most differences are not statistically significant among RNA-LLM, except for RiNALMo and ERNIE-RNA with respect to RNA-FM, RNABERT and RNAErnie. Details of statistical analysis in Supplementary Fig. S1a-d. e, Average $F_{1}$ per motif type on bpRNA-new. f, Performance accounting only non-canonical interactions on PDB-RNA.
@@
For a deeper analysis, we have extended the experiments on standard datasets having controlled levels of homology between training and testing partitions. Figure 2b shows the test $F_{1}$ for bpRNA dataset, with a partition at 80% sequence-identity cutoff between training (TR0+VL0) and testing (TS0) (see Methods). The results show that now, in a harder testing setup, most methods lowered performance in comparison to random partitions. Here, again, ERNIE-RNA and RiNALMo have both the best median scores in the test set, but in this case with a large drop in performance: $F_{1}=0.68$ and $F_{1}=0.64$, respectively. These two RNA-LLM are the only ones that achieve better performance than the classical prediction method ($F_{1}=0.60$). With median performance below it, RNA-FM has $F_{1}=0.52$, RNA-MSM has $F_{1}=0.42$, and RNAErnie has $F_{1}=0.41$. Both one-hot and RNABERT reached similar and very poor results, $F_{1}=0.35$ and $F_{1}=0.33$, respectively.
@@
Figure 2c shows the results of testing already trained models with the bpRNA-new dataset, which was designed to evaluate the prediction on new RNA families, never seen during training. The performance of all methods has been noticeably reduced, being the median of the RNA-LLM in all cases below the classical folding method ($F_{1}=0.71$). The only model with performance very close to the classical one is ERNIE-RNA, having a median $F_{1}=0.67$ (with large dispersion). In all the other cases, the median $F_{1}$ is below $0.45$. Additionally, note that the one-hot encoding moved up one more position in the ranking, slightly outperforming RNAErnie in this test set. This large dataset was used to analyze the prediction scores per structural motif, as seen in Fig. 2e. It can be seen that most models behave similarly for each motif type, with ERNIE-RNA showing a higher performance in stem base-pair predictions. Overall, ERNIE-RNA, RiNALMo and RNA-MSM achieved the best scores for each motif.
@@
Results on a more challenging dataset are shown in Fig. 2d, where every RNA-LLM was tested in a set of 62 RNA sequences whose secondary structures were inferred from high-resolution RNA X-ray 3D structures from the PDB dataset. Results clearly show a hard fall in results for all the RNA-LLM, with median performances below $F_{1}=0.40$, lower than the classical folding method, which achieved a moderate performance of $F_{1}=0.60$. Notably in this dataset, the one-hot encoding is within the top-3 methods, outperforming RNA-FM, RNAErnie and RNABERT; with almost the same performance that RNA-MSM, ERNIE-RNA and RiNALMo. It can be said that, for this prediction challenge, no RNA-LLM provides significant improvements of transfer learning over a one-hot encoding. This dataset also contains non-canonical interactions, although most of the methods fail to reach a median $F_{1}$ score higher than $0.2$ (Fig. 2f), with a slightly better performance for ERNIE-RNA and RiNALMo.
@@
For assessing inter-family performance, a family-fold cross validation in the ArchiveII dataset was performed. That is, one family was left out for testing per cross-validation fold, and the rest of the families were used for training. This eliminates most of the homology to the training set, providing a hard measure of performance and, thus, allowing estimating future performance on novel RNAs that do not belong to any known family. These results are shown in Fig. 3, along with the distributions of sequence lengths and minimum structural distances [38] between each test sequence and all the sequences in its corresponding training fold.

### v2

@@ Abstract
In recent years, inspired by the success of large language models (LLM) for DNA and proteins, several LLM for RNA have also been developed. These models take massive RNA datasets as inputs and learn, in a self-supervised way, how to represent each RNA base with a semantically rich numerical vector. This is done under the hypothesis that obtaining high-quality RNA representations can enhance data-costly downstream tasks, such as the fundamental RNA secondary structure prediction problem. However, existing RNA-LLM have not been evaluated for this task in a unified experimental setup. Since they are pre-trained models, assessment of their generalization capabilities on new structures is a crucial aspect. Nonetheless, this has been just partially addressed in literature. In this work we present a comprehensive experimental and comparative analysis of pre-trained RNA-LLM that have been recently proposed. We evaluate the use of these representations for the secondary structure prediction task with a common deep learning architecture. The RNA-LLM were assessed with increasing generalization difficulty on benchmark datasets. Results showed that two LLM clearly outperform the other models, and revealed significant challenges for generalization in low-homology scenarios. Moreover, in this study we provide curated benchmark datasets of increasing complexity and a unified experimental setup for this scientific endeavor.
Availability: The source code to reproduce all the experiments and results can be found in: https://github.com/sinc-lab/rna-llm-folding
@@ Introduction
Ribonucleic acid (RNA) plays a crucial role in many fundamental biological processes, such as gene expression, cell signaling, and post-transcriptional regulation [26, 72]. As in proteins, RNAs function and interaction with other molecules are deeply related to their structure. For example, determining RNA structure is essential for RNA-based therapeutics such as mRNA vaccines [42]. Among all RNA transcripts, only 5% is responsible for protein coding. At the same time, a very large remaining portion is non-coding RNA (ncRNA) [9], which in many cases adopt specific structures to perform important biological functions [8, 3]. Experimental results show that, to some degree, sequence determines the secondary structure, and thus, their function [40, 20, 67]. Therefore, the long-established and yet unresolved RNA secondary structure prediction is a major challenge in computational biology today [6, 7].
@@
RNA-LLM attempt to effectively embed RNA bases using deep representation learning, in particular the one developed in the field of natural language processing [24, 47]. Most RNA-LLM are based on bidirectional encoder representations from transformers (BERT), which was designed to generate context-sensitive distributed test token (word) representations [15]. The relevance of LLM for transfer learning from sequences to downstream tasks have been analyzed in detail for proteins [66, 31, 68]. In the case of RNA, the hypothesis is that LLM representation of nucleotide composition and sequence motifs can help characterize the structure and function of a sequence, analogously to how the meaning of a sentence is determined by the grammatical structure of natural language. Thus, word embedding techniques for natural language have been applied to bases for RNA sequences, obtaining a wide variety of BERT-based architectures trained on different databases. However, in spite that some theoretical reviews on the applications and utility of LLM in bioinformatics have recently appeared [32, 63, 43], to the best of our knowledge all the available RNA-LLM have not yet been comparatively evaluated for the secondary structure prediction task, considering a fair experimental setup, with the same datasets and structure prediction model. Since LLM are pre-trained models, it is important to analyze how they were trained to assess generalization capabilities on new sequences/structures, considering homology-aware partitions and cross-family predictions.
@@
Multiple sequence alignment (MSA) inspired by AlphaFold2 [27]. Pre-training with MLM and 20% random masking. Homology search and MSA of training sequences with RNAcmap3 database [12] as augmented input. Two modules: embedding and MSA transformer. Initial embedding layer and two learnable position-embedding layers. Stack of MSA transformer blocks with a residue and sequence attention layer followed by a feed-forward layer with GELU activation functions. Cross-entropy loss function.
@@
RNA large language models evaluated
In the last three years, a number of RNA-LLM have appeared in literature. The following RNA-LLM (summarized in Table 1) were included in this study according to their availability as open source tools.
RNABERT [1]: For the pre-training of the masked language modeling (MLM)[16] task in this LLM, $76,237$ human derived small ncRNAs from RNAcentral [62] were utilized. First, a token embedding randomly generates a 120-dimensional numerical vector that encodes the four RNA bases (A, C, G, U) and assigns the same vector to each base in the input RNA sequence. Second, the position embedding generates a 120-dimensional vector that encodes the position information of each base in the sequence. Third, the element-wise sum of token embedding and position embedding for each base in the input RNA sequence is fed to the Transformer layer. RNABERT architecture consists of a stack of 6 Transformers, each of which is composed of a multi-head self-attention mechanism followed by a feed-forward neural network. The weights of the last layer are trained by alternating between different tasks. The MLM task masks 15% of the bases randomly selected from the input RNA sequence and predicts the masked part using the neighboring bases, enabling a context sensitive embedding. The structural alignment learning (SAL) task is based on RNA structural alignment and aims to obtain closer embeddings for bases in the same column of reference alignment. The seed alignment for each family was downloaded from Rfam [28] as the reference structural alignment. The MLM task enables the position and context-sensitive embedding, and the SAL task enables the structural information embedding.
RNA-FM [11]: It is a 100 million parameter foundation model encoder based on the BERT original implementation and trained on 23.7 million unannotated ncRNAs from RNAcentral [62] database. Identical sequences were removed by applying CD-HIT [19] with a cut-off at 100%. The resulting dataset was used to train the foundation model in a self-supervised manner, reconstructing the input masked tokens as a pretext task. During pre-training under self-supervised training, around 15% of nucleotide tokens were randomly replaced with a special mask token. RNA-FM takes raw sequential tokens as input and embeds each nucleotide into a 640-dimensional vector. The architecture has 12 transformer encoder blocks as in BERT, which includes multi-head self-attention modules and feed-forward layers, with a final softmax layer to predict the output. The model was trained with MLM by predicting the original masked token with cross-entropy loss.
RNA-MSM [78]: Inspired by the success of AlphaFold2 [27] and the use of homologous sequences for the highly accurate prediction of protein structures, RNA-MSM is a multiple sequence alignment (MSA)-based RNA language model. It utilizes a set of homologous sequences that allows having a larger number of sequences for training. $4,069$ RNA families were downloaded from Rfam 14.7, totaling $3,087,138$ sequences. The RNAcmap3 database [12] for homolog search and sequence alignment was employed. Rfam families containing RNA sequences with experimentally determined structures were excluded to minimize potential overfitting for downstream tasks such as structural inference. This led to a total of $3,932$ Rfam families. Pre-training was based on MLM, with 20% random masking. The embedding module, with one initial embedding layer and two learnable position-embedding layers, encodes entries in the MSA separately. The architecture is made of a stack of MSA transformer blocks, each with a residue and sequence attention layer containing 12 heads with an embedding size of $768$, followed by a feed-forward layer, which summarizes approximately 96 million parameters. The use of structural information associated with the sequences was not included during the LLM pre-training. To provide a fair comparison with the other RNA-LLM, a single-sequence version of this method was used. This way, the embedding obtained corresponds only to the input sequence.
ERNIE-RNA [75]: This RNA pre-trained language model is based on the Enhanced Representation through Knowledge Integration (ERNIE) framework [60] and a modified BERT that incorporates base-pairing restrictions to be used with RNA. For training the model, a 34 million ncRNA dataset from the RNAcentral database was downloaded. After refining the vocabulary and removing redundant sequences, the final dataset consisted of 20.4 million RNA sequences. ERNIE-RNA was trained with MLM, which predicts the masked token with cross-entropy loss. The architecture has 12 transformer blocks, with 12 attention heads each. Every token in the input sequences is mapped to a $768$-dimensional vector, resulting in 86 million parameters. The main hypothesis is that ERNIE-RNA can learn functional and structural information thanks to the use of attention maps during pre-training.
RNAErnie [70]: It is also built upon the Enhanced Representation through Knowledge Integration (ERNIE) framework [60], together with multilayer and multihead transformer blocks. Pre-training was done with a corpus of approximately 23 million sequences extracted from the RNAcentral, and using self-supervised learning with multilevel random masking. The main difference with other works was the use of a motif-aware pre-training strategy involving motif-level and subsequence random masking, which can capture both subsequence and motif-level knowledge extracted from motif databases [30]. The architecture of RNAErnie shares the same architectural configuration as ERNIE 2.0 [61]: a 12-layer transformer with a hidden state embedding dimension of $768$. A block first tokenizes RNA bases in the sequence and subsequently feeds them into the transformer. Given the embeddings for every token in the RNA sequence, the RNAErnie basic block transforms the series of token embeddings into a $768\times L$ embedding using trainable parameters and then outputs the embedding of the RNA sequence. The total number of trainable parameters in RNAErnie is approximately 105 million.
RiNALMo [44]: It is the largest RNA language model to date with 650 million parameters pre-trained on 36 million unique non-coding RNA sequences from the RNAcentral [62] database augmented by Rfam [29], nt [52] and Ensembl [34]. To ensure diversity in each batch, the sequences were clustered with MMSeqs2 [57] into 17 million clusters, and then each batch contained a mixture of sequences sampled from different clusters. The architecture of RiNALMo is a BERT-style encoder-only Transformer [16]. Before passing to the Transformer, an RNA sequence is tokenized and represented as a $1280$-dimensional vector. RiNALMo consists of 33 Transformer blocks, where each block comprises a multi-head attention and a feed-forward network. The position of the tokens is encoded using rotary positional embedding [59]. Each multi-head attention has 20 heads. To improve pre-training efficiency, FlashAttention-2 [14] is employed. In the feed-forward network, two linear layers are used together with SwiGLU activation function [53]. Among the transformer modules, there are residual connections with layer normalization to stabilize training.
Materials and methods
Data
The following datasets have been used in this study. For all datasets, sequences longer than $512$ nucleotides were filtered to limit computational requirements [64]. That is, $512$ nt is the top length in each dataset.
ArchiveII dataset [56]: The most widely used benchmark dataset for RNA folding methods, containing RNA structures from 9 RNA families: 5s (ribosomal RNAs), srp (signal recognition particle), tRNA (transfer RNA), tmRNA (transfer messenger RNA), RNaseP (Ribonuclease P), grp1 (Glycine-rich RNA-binding protein 1), 16s (ribosomal RNA), telomerase and 23s (ribosomal RNA). The total number of sequences is $3,864$. This dataset is used in two training and testing split configurations. First, a random set of 5-fold partitions, following the original splits provided by the authors for the ArchiveII dataset [56]. We also perform a cross-family generalization analysis, training on all RNA families but one that is used as a test unseen family, and repeating for all families. That is, a leave-one-family-out strategy.
bpRNA dataset [54]: The same train and test sets as used in SPOT-RNA[54]. It is a non-redundant set of RNA sequences at 80% sequence-identity cutoff with CD-HIT-EST, with annotated secondary structure from bpRNA34 [19]. This filtered dataset of $13,419$ RNAs is randomly divided into $10,814$ RNAs for training (TR0), $1,300$ for validation (VL0), and $1,305$ for an independent test (TS0). In this dataset, each model was trained with TR0+VL0 and tested with TS0.
bpRNA-new dataset [54]: This dataset was derived from Rfam 14.2 [29], containing new RNA families different from the bpRNA dataset. This test dataset has $5,401$ sequences and was built to assess cross-family model generalization. This dataset is used as an additional test set for the models trained with the bpRNA dataset.
PDB-RNA dataset [50] The PDB sets TR1 (training), VL1 (validation) and TS1 (testing) are the same sets as in SPOT-RNA, prepared by downloading all the high-resolution ($<3.5$A°) RNA X-ray structures from the PDB dataset on March 2, 2019. The numbers of structures for TR1, VL1 and TS1 are $120$, $30$ and $62$, respectively, after removing homologous sequences between and within the sets by CD-HIT-EST at the lowest allowed sequence identity cut-off of 80%. In this dataset each model was trained with TR1+VL1 and tested with TS1. This small training set is used to test the LLM capabilities in a very hard setup for any transfer learning approach.
Self-supervised RNA-LLM and prediction model
For the comparative analysis, embeddings of each RNA-LLM feed the same deep learning (DL) architecture (Figure 1a-e) for RNA secondary structure prediction. The rationale behind this is that it should not be necessary to optimize the classifier hyperparameters for each LLM, since those have complex architectures and have already been trained with a large amount of data. Thus, according to the transfer learning paradigm, the embeddings should have enough information to solve the downstream task. Moreover, any difference measured in performance will be due to the LLM and not because of the classifier architecture. This is the only trainable part of the experimental setup, which uses the train-test partitions defined for each benchmark dataset.
Sequence one-hot embedding was defined as usual using 4 positions for A, U, G and C (Fig. 1b). Each LLM was used with the pre-trained weights provided, according to instructions in the official repositories. Models were frozen, that is, not re-trained nor fine-tuned. Per-nucleotide embeddings were extracted for each sequence in each benchmark dataset, obtaining a $d\times L$ tensor, with $d$ the embedding dimension and $L$ the sequence length. The embeddings feed the secondary structure prediction network described in Figure 1a. This architecture was designed following the one used in RiNALMo [44], RNA-FM [11], RNA-MSM [78], and ERNIE-RNA [75]. A fully connected layer reduces the input dimension to $M/2$ (being $M\ll d$), in order to obtain the same dimension for all RNA-LLM. An outer concatenation approach was used to transform the $M/2\times L$ projection to a $M\times L\times L$. The $M$ dimension can be interpreted as channels of a $L\times L$ image. Then, this tensor passes through two 2D ResNet blocks [23] with the following configuration: a first ResNet with convolutions of kernel size 1 and a second one with kernel size 3. Both ResNet blocks include instance Normalization and ReLU activation function. The same DL architecture was trained and tested using each RNA-LLM and dataset. For training, Adam optimizer was used with a learning rate of $0.0001$, a batch size of 4 and binary cross-entropy as the loss function. A fixed number of 15 epochs was used in order that all RNA-LLM have equal possibilities for training. The number of epochs and the learning rate were determined in preliminary experiments, ensuring that there was no overfitting nor underfitting of the secondary structure prediction model for any of the RNA-LLM embeddings used as input. Each dataset was used with its own training and testing partitions, except for bpRNA-new, which was only used as a testing partition for the prediction model already trained on bpRNA. All experiments were run on 3 servers, each with 2 NVIDIA RTX A5000 GPU.
Figure 1: RNA-LLM embeddings and deep neural architecture of the prediction model. a, Flow diagram for RNA secondary structure predictions. Each LLM was downloaded from its official repository, frozen and used to get per-nucleotide embeddings for each sequence in the benchmark datasets. The embeddings go through a fully connected layer, outer concatenation, two 2D ResNet blocks and a final 2D convolution. This flow is explained in detail in aligned panels b-e. b, One-hot encoding, of size $4\times L$, and LLM representation with size $d\times L$, where $d$ is the embedding dimension and $L$ is the padded sequence length. A fully connected layer reduces the input dimension to $M/2$. c, The $M/2\times L$ projection is transformed to a $M\times L\times L$ tensor using outer concatenation. The position $(i,j)$ contains the concatenated representation of nucleotides $i$ and $j$. d, Then, there are two 2D ResNet blocks with 2D Convolution, Instance Normalization and ReLU. ResNet Blocks 1 has kernel size 1 and ResNet Blocks 2 has kernel size 3. e, A final convolution yields the output scores as a $L\times L$ connection matrix. f-l, RNA-LLM embeddings projected with UMAP for dimensionality reduction. The RNA families of the ArchiveII dataset are highlighted with different colors. Sequences within the same RNA family are expected to be close in the dimensionally reduced space.
For the comparison of RNA-LLM based predictors with baseline methods, the most widely-used pure-thermodynamic structure predictors were included: RNAstructure [46], RNAfold [33] and LinearPartition-V [76]. We have also incorporated hybrid methods, that is, a mix of thermodynamic and machine learning approaches, such as LinearPartition-C [76] and MXFold2 [50]. Finally, we also included the most recent pure DL prediction methods: UFold [18], REDfold [10], and sincFold [7].
Performance measures
Results are reported with the following performance measures.
Base pairs metrics: The focus of performance measures is on the predicted base pairs in comparison to a reference structure [35]. Pairs that are both in the prediction and the reference structure are true positives (TP), while pairs predicted but not in the true structure are false positives (FP). Similarly, a pair in the reference structure that is not predicted is a false negative (FN), and a pair that is neither predicted nor in the true structure is a true negative (TN). A widely used metric is the recall (or sensitivity), defined as the ratio of TP to all the true pairs (TP+FN). It is a measure of how many predicted pairs are true. Insted, the precision is defined as the ratio of TP to all the predicted pairs (TP+FP). This relation between TP and FP is very important in the context of class imbalance, because FP can be a large number in comparison to TP, and this is not reflected in the recall. The $F_{1}$ score is the harmonic mean of precision and recall, thus summarizing both measures of performance in a single value. Therefore, in this work the $F_{1}$ score is used as the global performance measure for the comparison of the different methods. It is defined as $F_{1}=\frac{2TP}{2TP+FP+FN}$.
Structural motifs: Stems, multiloops, internal loops, bulges, hairpin loops, dangling ends, and external loops were extracted using the bpRNA toolkit [13] on the reference dot-bracket structures. All reference and predicted base-pairs were extracted for positions of each motif type, and average $F_{1}$ computed separately.
Structural similarity: to complement the $F_{1}$ score as an evaluation metric, we have also calculated a structural measure, the Weisfeiler-Lehman graph kernel (WL) metric [48]. The WL metric first assigns to each node (nucleotide) in the graph (secondary structure) a label representing its local structural information. Then, a label propagation step iterates over the nodes and updates their labels based on the labels of their neighboring nodes. Finally, a hash function is computed that aggregates these labels to generate a feature vector. The WL is defined as inner product $WL(G_{1},G_{2})=\Phi(G_{1})\cdot\Phi(G_{2})$, where $\Phi(G_{i})$ represents the feature vector of graph $G_{i}$ obtained by aggregating the labels through the hash function.
@@ Results
Figure 2 shows the comparative results among the RNA-LLM here reviewed in terms of $F_{1}$ violin plots. In each plot the results are presented in order, from best (left) to worst (right) median $F_{1}$. The best thermodynamic and DL-based predictors were used as baselines in the plots (details in Supplementary Table 1).
@@
Figure 2a shows the results for 5-fold random partitions on the ArchiveII dataset. It can be seen that most predictors achieve a median performance above $F_{1}=0.60$, except only for the one based on one-hot encoding ($F_{1}=0.57$). The top-3 RNA-LLM achieve a very high performance ($F_{1}>0.90$). Those are ERNIE-RNA and RiNALMo, both with $F_{1}=0.95$, and RNA-FM with $F_{1}=0.91$. RNAErnie and RNA-MSM achieve an intermediate performance, $F_{1}=0.76$ and $F_{1}=0.74$, respectively. Finally, RNABERT is the RNA-LLM with the lowest average performance ($F_{1}=0.62$). For this random $k$-folding, all RNA-LLM achieve higher performance than the classical prediction method ($F_{1}=0.61$). Remarkably here, the DL-based method achieves the highest median performance ($F_{1}=0.97$). This result shows that the neural architecture used in the classifier has the capability of obtaining high performance in predictions. However, it is well-known that the sequence homology between train and test at random partitions favors methods based on deep learning; thus performance tends to be overly optimistic.
@@
Figure 2: Comparative results among RNA-LLM on the RNA secondary structure prediction task for different benchmark datasets of increasing complexity. Each method has a different color. The thermodynamic prediction method (LinearPartition-V, dashed black) and the DL-based prediction method (sincFold, solid black line) are added as baselines. Next to RNA-LLM performances, sequence length distribution for each dataset is shown in blue. a, ArchiveII 5-fold random cross-validation. A Wilcoxon test for paired samples with Bonferroni correction indicates that all differences are statistically significant ($P<0.0001$, $N=3,864$). b, bpRNA train-test partitions with controlled homology. All differences are statistically significant ($P<0.0001$, $N=1,305$) except for one-hot and RNABERT. c, bpRNA-new dataset, for RNA families not seen during training. All differences are statistically significant ($P<0.0001$, except for RNAErnie versus RNA-MSM and one-hot with $P<0.05$, $N=5,401$). d, PDB-RNA dataset, with RNA sequences extracted from PDB. Most differences are not statistically significant among RNA-LLM, except for RiNALMo and ERNIE-RNA with respect to RNA-FM, RNABERT and RNAErnie. Details of statistical analysis in Supplementary Fig. S1a-d. e, Average $F_{1}$ per motif type on bpRNA-new. f, Performance accounting only non-canonical interactions on PDB-RNA.
For a deeper analysis, we have extended the experiments on standard datasets having controlled levels of homology between training and testing partitions. Figure 2b shows the test $F_{1}$ for bpRNA dataset, with a partition at 80% sequence-identity cutoff between training (TR0+VL0) and testing (TS0).
@@
The results show that now, in a harder testing setup, most methods lowered performance in comparison to random partitions. Here, again, ERNIE-RNA and RiNALMo have both the best median scores in the test set, but in this case with a large drop in performance: $F_{1}=0.68$ and $F_{1}=0.64$, respectively. These two RNA-LLM are the only ones that achieve better performance than the classical prediction method ($F_{1}=0.53$) and very close performance to the DL-based method ($F_{1}=0.68$). With median performance below it, RNA-FM has $F_{1}=0.52$, RNA-MSM has $F_{1}=0.42$, and RNAErnie has $F_{1}=0.41$. Both one-hot and RNABERT reached similar and very poor results, $F_{1}=0.35$ and $F_{1}=0.33$, respectively.
@@
Figure 2c shows the results of testing already trained models with the bpRNA-new dataset, which was designed to evaluate the prediction on new RNA families, never seen during training. The performance of all methods has been noticeably reduced, being the medians of almost all RNA-LLM below the classical folding method ($F_{1}=0.67$). The only RNA-LLM with the same performance as the classical one is ERNIE-RNA. In all the other cases, the median $F_{1}$ is below $0.45$. RiNALMo achieved the same median performance as the DL-based baseline ($F_{1}=0.44$). Additionally, note that the one-hot encoding moved up one more position in the ranking, slightly outperforming RNAErnie in this test set. This large dataset was used to analyze the prediction scores per structural motif, as seen in Figure 2e. It can be seen that most models behave similarly for each motif type, with ERNIE-RNA showing a higher performance in stem base-pair predictions. Overall, ERNIE-RNA, RiNALMo and RNA-MSM achieved the best scores for each motif.
@@
Results on a more challenging dataset, closer to a real-world application, are shown in Figure 2d, where every RNA-LLM was tested in a set of 62 RNA sequences whose secondary structures were inferred from high-resolution RNA X-ray 3D structures from the PDB dataset. Results clearly show a hard fall in results for all the RNA-LLM, with median performances below $F_{1}=0.40$, lower than the classical folding method ($F_{1}=0.67$) and the DL-based method ($F_{1}=0.58$). Notably in this dataset, the one-hot encoding is within the top-3 methods, outperforming RNA-FM, RNAErnie and RNABERT; with almost the same performance that RNA-MSM, ERNIE-RNA and RiNALMo. It can be said that, for this prediction challenge, no RNA-LLM provides significant improvements of transfer learning over a one-hot encoding. This dataset also contains non-canonical interactions, although most of the methods fail to reach a median $F_{1}$ score higher than $0.2$ (Fig. 2f), with a slightly better performance for ERNIE-RNA and RiNALMo.
Additionally to the $F_{1}$ score, we have also calculated for all the datasets analyzed the WL metric, which was proposed to capture structural information between predictions and the corresponding references. Supplementary Table 1 shows that the performances have the same trend to those results reported and analyzed with $F_{1}$.